find the equation in the form of y=mx+b through (3,-5) parallel to y=4x+3
Answer by Maths68(1474)

Given
Point (x, y)=(3,-5)
Line:
y=4x+3
Compare above equation with the equation of line slope-intercept form
y=mx+b
y=(4)x+3
m=4 and b=3
Slope of the given line m = 4 and y-intercept = b = 3
Since given line is parallel to the required line; slopes of the both line will be same.
================================================================
Now we have a point (3,-5)) and slope (4) of the required line we can easily find the required line put these values in the equation of slope-intercept form to find the y-intercept of the required line
y=mx+b
-5=(4)(3)+b
-5=12+b
-5-12=b
-17=b
b=-17
y-intercept of the required line =b=-17
================================================================
Put the values of ‘m’ and ‘b’ in equation of the line slope-intercept form
y=mx+b
y=(4)x+(-17)
y=4x-17
Above equation is the required equation of the line in slope-intercept form